OpenStudy (theeric):

Entropy Question: How can water be heated from 273K to 363K while having entropy change of the entire system be zero?

OpenStudy (theeric):

The problems that we've been working on have involved entropy change as water is heated by heat baths or copper... My confusion is that the entropy change of a system being zero means that the process is reversible, and I don't see how heating water could be reversible. I think the process might be reversible if something else in the system is made cold in the process, but that can't happen on it's own. So work would need to be done, maybe? These ideas are not completely solid to me. I have a professor who is very disorganized on the blackboard, and this professor writes physics math faster than I can understand it sometimes. Thanks for any help!

OpenStudy (isaiah.feynman):

By entropy you mean heat transfer? Cos that what I'm thinking.

OpenStudy (theeric):

Entropy is associated with energy exchange and temperature. I don't know a whole lot about it, but it's more than transferred energy. From Wikipedia: \(\Delta S=\huge{\int}\normalsize\dfrac{\rm dQ}{T}\)

OpenStudy (isaiah.feynman):

Haven't studied this yet.

OpenStudy (theeric):

I got ya. I'm just studying it now! Thanks for looking!

OpenStudy (theeric):

\(Q\) is transferred energy, by the way!

OpenStudy (isaiah.feynman):

Then dQ means what? Change in transferred energy?

OpenStudy (isaiah.feynman):

@theEric

OpenStudy (isaiah.feynman):

I'm interested in this.

OpenStudy (theeric):

I don't know much about it! But yeah, that's the infinitesimal change in energy. Here's some things on entropy: http://hyperphysics.phy-astr.gsu.edu/hbase/thermo/seclaw.html#c4

OpenStudy (theeric):

The first law of thermodynamics is that energy is conserved. So \(\Delta E=Q+W\) where \(W\) is work done on the system and \(Q\) is energy transferred to the system. But that alone doesn't explain why things happen the way they do. For example, water doesn't spontaneously become hot by taking energy from around it. This would be fine with conservation of energy, but it doesn't happen. And so the second law of thermodynamics comes into play.

OpenStudy (isaiah.feynman):

W should be the work done by the system since its positive.

OpenStudy (isaiah.feynman):

If the entropy is zero. Then the system must be in thermal equilibrium. Energy transfer should be zero in this case.

OpenStudy (theeric):

So, I do have the equation and descriptions right. It makes sense, though. If the system does work, it loses that energy. If work is done ON the system, you see a positive change in energy of the system.

OpenStudy (theeric):

Energy can be transferred to the system even with thermal equilibrium. Example: We have boiling water in a closed system. When the water is heated, more water vapor will escape but maintain the boiling temperature until all of the water has evaporated.

OpenStudy (theeric):

Tricky stuff, huh?

OpenStudy (theeric):

That situation is somewhat idealized, of course...

OpenStudy (theeric):

If the entropy change is zero, though, it means that the process is reversible. The process can go either way.

OpenStudy (theeric):

Part of the law holds that \(\Delta S\ge 0\) always.

OpenStudy (theeric):

Many processes have positive entropy.. Like heating water... as far as I know.

OpenStudy (theeric):

I'm gonna log out and get ready to sleep. Have a good night! Have fun learning about thermodynamics if you choose to!

OpenStudy (theeric):

Take care! :)

OpenStudy (isaiah.feynman):

You TOO sir.

OpenStudy (theeric):

I think the answer might have been using an infinite number of heat baths. Either way, since entropy was constant through every quasistatic process, each process is considered to be quasistatic and adiabetic.